# PatchTrack: A Comprehensive Analysis of ChatGPT’s Influence on Pull Request Outcomes

### v1

@@ Abstract
The rapid adoption of large language models (LLMs) like ChatGPT in software development has introduced new ways for developers to interact with AI, particularly in pull request workflows. While prior research has examined AI-generated code quality, there is limited understanding of how ChatGPT is utilized in real-world pull request decision-making and how its suggestions influence patch integration and rejection. To explore these aspects, we analyze self-admitted ChatGPT usage (SACU), where developers explicitly disclose their reliance on ChatGPT within pull request discussions. Our study examines 338 pull requests (285 merged, 53 closed) across 255 GitHub repositories, containing 645 ChatGPT-generated code snippets and 3,486 patches. We introduce PatchTrack, a classification tool that determines whether ChatGPT-generated patches were applied (PA, 115 cases), not applied (PN, 64 cases), or not suggested (NE, 106 cases). Our findings reveal that full adoption of ChatGPT-generated code is rare, developers frequently modify or selectively integrate AI-generated patches to align with project constraints, with a median integration rate of 25%. Through qualitative analysis, we identify key factors influencing patch integration and pull request rejection, including scope misalignment, maintainability concerns, redundant solutions, and procedural barriers such as incomplete documentation or administrative policies. By providing empirical insights into ChatGPT’s role in pull request workflows, this study informs developers, maintainers, and educators on the evolving use of generative AI in collaborative software development. It also lays the groundwork for future research on optimizing AI-assisted development, improving transparency in AI adoption, and enhancing patch integration workflows.
@@
AI-assisted workflows are increasingly shaping software development, with tools such as ChatGPT and GitHub Copilot that improve automation, developer productivity, and collaboration [1, 2, 3, 4, 5, 6, 7]. Prior studies have explored their role in code generation, debugging, and documentation [8, 9, 10, 11], yet their impact on collaborative workflows, particularly in pull request decision-making, remains less understood. Understanding how developers integrate, modify, or reject AI-generated code in pull requests is essential to improve AI-assisted collaboration.
Pull requests are central to modern software engineering, enabling collaborative contributions and rigorous code review processes [12, 13, 14]. Despite their importance, PR workflows remain effort-intensive, requiring developers to evaluate code quality, resolve integration conflicts, and ensure alignment with project goals [6, 7]. The long-standing vision of the Programmer’s Apprentice [15] envisioned AI as a reasoning collaborator, offering suggestions, help with debugging and design insights, while leaving key decisions to developers. Recent advances in generative AI have brought this vision closer to reality, with tools now increasingly integrated into everyday development workflows. However, the effectiveness of such tools in practical pull request scenarios - especially to influence integration, revision, or rejection decisions - remains largely unclear [16, 17, 18, 19].
Although prior research has explored predictive models and recommendation systems for pull request decision making [20, 21, 22], AI-assisted software engineering research has focused primarily on code quality, debugging, and documentation improvements[23, 24, 25, 26]. Studies evaluating GitHub Copilot and ChatGPT highlight their productivity benefits[27, 28, 29, 30, 31, 32] but offer limited insights into collaborative pull request workflows. Furthermore, while ChatGPT-generated code has been studied in the context of integration rates and modification trends [33, 34], the decision-making processes behind patch acceptance, modification, and rejection remain largely unexplored.
This study addresses these gaps by analyzing ChatGPT conversations shared in merged and closed pull requests. We classify AI-assisted pull request outcomes and examine the factors that influence developer decisions to accept, modify, or reject ChatGPT-generated patches. Unlike previous work that focused on AI-generated code retention [33] or security concerns [34], we investigated the broader role of ChatGPT in pull request decision-making, including cases where pull requests are closed despite AI involvement. Our study builds on research by Champa et al. [18] and Chouchen et al. [35], situating ChatGPT’s role within real-world pull request workflows to assess its impact on collaboration, decision-making, and code review efficiency.
To examine ChatGPT’s influence on pull request workflows, we analyze both merged and closed pull requests where developers shared ChatGPT-generated content, including both code patches and conceptual guidance, leveraging the platform’s link-sharing feature introduced in May 2023 [36]. Unlike studies that primarily assess AI-generated code quality or developer modifications [33, 34], our study investigates ChatGPT’s broader role in pull request decision-making, including its impact on pull request closures. This includes cases where ChatGPT provides direct code suggestions, as well as instances where it offers conceptual guidance, debugging strategies, or alternative solutions. Building on work by Champa et al. [18] and Chouchen et al. [35] exploring developer interactions with ChatGPT, we examine its potential to streamline decision-making, enhance collaboration, and improve code review processes. By classifying pull request outcomes, our analysis provides empirical insight into ChatGPT’s strengths and limitations in collaborative software development, moving beyond controlled experiments to real-world workflows.
To support this investigation, we extend the DevGPT dataset[37], incorporating additional repositories across diverse programming languages and project contexts. This research expands on an initial study presented at ASE 2024[38], where we introduced PatchTrack and conducted a preliminary analysis of AI-generated patches. Our study significantly extends this work by (1) improving PatchTrack’s classification algorithm, improving the precision of identifying ChatGPT’s influence on pull request outcomes, (2) expanding the dataset to include more closed pull requests, allowing a deeper analysis of AI-assisted pull request rejections, and (3) broadening the scope of research beyond patch application to investigate the reasons pull requests are closed despite AI assistance, such as scope misalignment, patch quality and project constraints. These advancements provide a comprehensive assessment of the role of ChatGPT in pull request workflows, offering insight into the larger impact of AI on collaborative software development and decision-making.
This study provides a foundation for future research in AI-assisted software development by empirically analyzing ChatGPT-assisted pull request workflows, identifying key factors influencing patch integration and rejection, refining automated classification methods, and offering practical recommendations for responsible AI adoption in collaborative coding environments. Our key contributions are as follows.
We analyze ChatGPT-assisted pull request workflows, classifying outcomes to assess AI’s influence on patch integration and rejection, helping developers optimize AI-assisted development.
We enhance PatchTrack’s classification accuracy through benchmark testing and manual validation, improving patch tracking reliability.
We expand the dataset with additional repositories and closed pull requests, enabling a broader analysis of ChatGPT’s role in pull request decisions and guiding AI adoption strategies.
We perform a qualitative analysis of pull request decision-making, identify key factors that affect the integration of AI patches, and provide actionable insights for developers.
We offer practical recommendations for developers, researchers, and educators on responsible AI integration in software engineering workflows.
@@
We release a replication package [39] with datasets, classification tools, and scripts to support reproducibility and future research.
@@ 2.2 Classifying Developer Interaction with ChatGPT
Figure 1 presents a typical GitHub pull request workflow augmented with ChatGPT interactions, which can originate from the pull request author or a reviewer. Consider a scenario in which a developer, intending to fix a bug, add a feature, or improve an existing one, initiates a change to a file named Foo. Following a pull-based development workflow [41], the developer creates a new branch, commits changes, and eventually merges the pull request into the main repository of the project.
@@
If the code snippets are not applied, or modified beyond recognition (resulting in a semantic clone), Foo may remain unchanged or include changes influenced by the discussed concepts. This scenario is defined as “Patch Not Applied (PN)”.
@@ 3.1 Research Questions
RQ1: How are ChatGPT-generated patches (PA, PN, and NE) distributed and utilized in merged pull requests under the self-admitted ChatGPT usage? This RQ investigates the distribution of ChatGPT-generated patches in merged pull requests and explores how developers incorporate them into their workflows. It classifies pull requests into three categories: (i) patches that were applied (PA), (ii) patches that were suggested but not applied (PN), and (iii) cases where ChatGPT did not generate patches (NE). This classification provides insights into the extent of AI-generated patch adoption and its influence on development practices.
@@ 3.2.1 Step 1: Identification of merged pull requests containing ChatGPT conversation shared links:
This expanded search resulted in the identification of 187 new merged pull requests and 14 closed and unmerged pull requests that met our criteria. After filtering out inactive or ‘toy’ projects and verifying link functionality, our updated dataset now includes a total of 338 pull requests from 255 unique repositories, providing a robust foundation for analyzing the integration of ChatGPT conversations in software development workflows. As shown in Figure 4, we analyzed repository metrics such as stars, forks, commits, watchers, and language diversity to assess project maturity. The repositories varied significantly in size and engagement. The stars ranged from 0 to 81k (median: 58), and the forks ranged from 0 to 36,655 (median: 31), indicating both niche and widely recognized projects. The commits also showed a wide variation, with some repositories exceeding 36k. Watchers, a measure of visibility, were highly skewed, peaking at over 1.2 million (median: 946), reflecting differences in community interest and activity. These repository attributes have been used in previous studies to measure project diversity and participation [45, 46, 47].
The dataset includes repositories written in 123 distinct programming languages, demonstrating a high level of linguistic diversity. The most frequently used languages are Shell (85.5%), JavaScript (82.0%), HTML (62.0%), TypeScript (50.0%), Python (54.1%), and CSS (54.0%), indicating a strong presence of web development, scripting, and data processing projects. Other widely used languages include Dockerfile (49.0%), Makefile (33.0%), SCSS (22.4%), Java (19.0%), C (18.4%), and C++ (14.5%), reflecting a mix of infrastructure, compiled languages, and styling frameworks.
Although this extensive language diversity suggests that the dataset covers a broad spectrum of software development domains, its distribution is not fully representative of GitHub’s overall ecosystem. Compared to GitHub’s reported trends [48], our dataset overrepresents Shell scripting, web development technologies (HTML, CSS, JavaScript, TypeScript), and infrastructure-related files (Dockerfiles, Makefiles), while underrepresenting widely used languages such as Python and Java. According to GitHub’s Octoverse 2024 report [48], Python has become the most widely used programming language, surpassing JavaScript, and Java, C and C++ also rank highly in the use of the repository. This discrepancy suggests that developers may be more inclined to use ChatGPT for routine scripting, configuration management, and front-end development, where AI-generated suggestions align with established patterns and automation-friendly tasks. In contrast, in domains requiring complex reasoning, such as data science or large-scale backend engineering, developers can rely on specialized frameworks and tools beyond ChatGPT’s current capabilities.
Consequently, while our findings provide strong insights into ChatGPT’s role in pull request workflows within scripting and web development contexts, further research is needed to determine whether similar patterns hold in other software engineering domains. Future work could mitigate this bias by sampling repositories based on GitHub language distribution and investigating AI’s role in diverse development environments, including scientific computing, embedded systems and enterprise software.
@@ 3.2.2 Step 2: Extracting pull requests and ChatGPT patches
The pull request and ChatGPT patches are central for analysis in RQ1. Our dataset consists of links to pull requests, ChatGPT links, and patches (if any) extracted from the conversation. The process of extracting patches from ChatGPT conversations was not completely automated, especially in the extended dataset. We noted that OpenAI’s updated terms of service (as of November 14, 2023) no longer permit users to automatically extract data from conversation outputs [49].
To overcome this challenge, we manually converted all the shared links obtained in Step 1 into HTML format and performed scraping on the local files without making HTTP requests to the actual links. pull request patches are extracted using the GitHub API /pulls/ endpoint and saved as “diff” files. At the end of this step, we have two sets of patches for each pull request, namely ChatGPT and pull request patches. This data serves as input for the patch classification step.
@@ 3.2.3 Step 3: Patch classification
In this step, and the subsequent two steps, we utilize the PatchTrack patch detection tool, developed by adapting PaReco [50]. While PaReco is designed for the detection of code clones within divergent variant fork software families–identifying missed opportunities, effort duplication, and split cases between mainline and variant branches–it is not directly applicable for detecting ChatGPT patches. In contrast, PatchTrack specifically identifies ChatGPT-shared links, extracts patches, and analyzes prompts. Although PaReco supports exact token matching, its reliance on a Bloom filter data structure leads to higher false positive rates, which can undermine patch identification precision. PatchTrack addresses this limitation by using a temporary array of hash tables for token matching, allowing precise matches with reduced false positives. Hash tables are particularly suitable for exact data retrieval, offering an average-case time complexity of for lookups [51, 52], which contrasts favorably with the lookup time associated with Bloom filters. This design significantly improves PatchTrack’s efficiency and accuracy, making it more effective for exact token matching in diverse software development contexts.
To classify a patch suggested by ChatGPT for a specific pull request, PatchTrack compares it with all other patches in the pull request. PatchTrack performs normalization by removing whitespace, comments, converting characters to lowercase, removing non-ASCII characters, and applying other optimizations tailored to the file types identified by the file extensions of each file. Importantly, these normalization steps are applied to both the ChatGPT and pull request patches being classified. Currently, PatchTrack supports 38 different file types, primarily based on the dataset utilized in this study. These encompass the most commonly used file types within our collected data, enabling the tool to recognize and adapt to the unique syntax and comment styles of each programming language. The architecture of PatchTrack is designed with scalability in mind, allowing easy extension as needs arise or the dataset evolves, ensuring accurate and adaptable patch classification.
Building on methodologies established by tools such as PaReco [50] and ReDeBug [53], PatchTrack employ n-gram tokenization to facilitate precise patch detection. While PaReco and ReDeBug typically utilizes an n-gram value of , our tool uses to accommodate the often smaller size of ChatGPT patches. A detailed comparison of various -gram values is presented in Section 4.1, where we discuss our findings that validate the choice of as optimal for this study.
@@
A patch is classified as PA, PN, NE, “cannot classify” (CC), or “error encountered” (EE). This classification occurs at the Hunk level of each pull request patch. For each hunk, PatchTrack compares the n-gram tokens of the added lines (lines prefixed with ‘’) with those of the ChatGPT codesnippet. If at least one match is found, the patch is classified as PA; otherwise, it is classified as PN. In the absence of a ChatGPT code snippet for a given pull request case, all patches are classified as NE. Currently, PatchTrack supports all file types present in our dataset. Should a file type not yet be supported, it would be classified as CC, indicating areas for future enhancement rather than a limitation. Furthermore, any processing errors encountered during patch analysis lead to an EE classification, typically arising from issues within the tool’s processing functions.

### v2

@@ Abstract
The rapid adoption of large language models (LLMs) like ChatGPT has introduced new dynamics in software development, particularly within pull request workflows. While prior research has examined the quality of AI-generated code, little is known about how developers actually use these suggestions in real-world collaboration. We analyze 338 pull requests from 255 GitHub repositories containing self-admitted ChatGPT usage, including 645 AI-generated snippets and 3,486 developer-authored patches. We introduce PatchTrack, a tool that classifies whether ChatGPT patches were applied, not applied, or not suggested, enabling fine-grained analysis of AI-assisted decisions. Full adoption of ChatGPT code is rare: the median integration rate was 25%. A qualitative analysis of 89 pull requests with integrated patches revealed recurring patterns of structural integration, selective extraction, and iterative refinement, showing that developers typically treat ChatGPT’s output as a starting point rather than a final implementation. Even when code was not directly adopted, ChatGPT influenced workflows through conceptual guidance, documentation, and debugging strategies. Integration decisions were shaped by scope, architectural fit, contributor role, and review norms. This study offers empirical insight into how generative AI is used in collaborative software development, showing that its impact extends beyond patch generation to broader decision-making. Our findings inform the design of AI-assisted tools, clarify patch adoption behavior, and support more transparent and effective use of LLMs in practice.
@@
AI-assisted workflows are increasingly shaping software development, with tools such as ChatGPT and GitHub Copilot that promise improved automation, productivity, and collaboration [1, 2, 3, 4, 5]. Prior work has examined their utility in code generation, debugging, and documentation [6, 7, 8, 9], but most studies have focused on controlled settings or isolated tasks. Their impact on collaborative workflows, particularly pull request (PR) decision-making, remains poorly understood [10, 11]. Understanding how developers integrate, modify, or reject AI-generated code in real-world PRs is essential to improving AI-assisted collaboration.
Pull requests are central to modern software engineering, enabling peer review, discussion, and integration [12, 13, 14]. Traditional studies have shown that PR outcomes depend on factors like patch size, contributor history, and code quality [15, 16, 17, 18, 19]. Recently, tools like GitHub Copilot have entered the PR workflow [20], offering autogenerated summaries and suggestions to speed up reviews. These advances echo the long-standing vision of the Programmer’s Apprentice [21], in which AI augments rather than replaces human reasoning. Yet, the effectiveness of such tools in shaping real pull request decisions remains poorly understood.
Although prior research has explored predictive models and recommendation systems to support pull request decision-making [22, 23, 24], these studies typically rely on social and technical metadata, such as contributor history or patch size, rather than on the content of patches themselves. In contrast, research in AI-assisted software engineering has primarily focused on improving code quality, debugging, and documentation through generative models [25, 26, 27, 28]. While these approaches demonstrate technical utility, they do not investigate how AI-generated outputs are judged, adapted, or discarded in collaborative workflows such as pull requests.
Recent studies evaluating GitHub Copilot and ChatGPT emphasize productivity benefits, showing that LLMs can accelerate code writing and reduce developer effort [29, 30, 31, 32, 33, 34]. However, these evaluations are often limited to controlled settings or task-specific metrics, and they rarely examine how AI-generated content is evaluated or integrated in collaborative workflows. Even studies of ChatGPT-generated code that report on integration rates and patch retention [10, 11] do not capture the nuanced decisions developers make during pull request discussions.
Prior studies on pull request decision-making have primarily focused on whether patches are accepted or rejected, often examining factors like patch size, review effort, and contributor history [15, 16, 17, 18]. To move beyond binary merge decisions, we propose a more nuanced view of developer and AI interaction. Rather than asking whether a patch is accepted, we examine whether developers apply ChatGPT-suggested code directly, modify it significantly, act only on conceptual guidance, or ultimately reject it. These distinctions offer richer insight into how AI suggestions shape developer decisions, even when the code itself is not integrated.
Meanwhile, developers are increasingly incorporating ChatGPT into daily development [35], yet recent evaluations of automated code review tools show that such systems often overlook human decision-making context, leading to challenges in accuracy, trust, and integration [36]. Hallucinated or incorrect code further undermines adoption, as developers must verify or revise suggestions [37]. Multi-hunk patches pose additional challenges for large language models due to limitations in semantic coordination and global reasoning [38]. While these technical and socio-technical barriers are increasingly recognized, few studies have examined them empirically in the context of real-world pull request outcomes.
This study addresses that gap by analyzing 338 pull request conversations from open-source projects involving self-admitted ChatGPT usage (SACU). We introduce PatchTrack [39], a tool designed to classify patch outcomes and usage patterns under AI assistance. Originally presented as a poster at ASE 2024, we significantly extend PatchTrack’s capabilities by improving classification precision, validating its outputs, and applying it to a larger, SACU-enhanced version of the DevGPT dataset [40]. This enables a scalable and systematic analysis of how ChatGPT-generated suggestions are integrated, adapted, or rejected in collaborative workflows.
Unlike prior work that focuses primarily on integration rates or isolated productivity metrics [29, 30, 31, 32, 33, 34, 41], we adopt a behavioral lens that examines how developers interpret, adapt, or disregard ChatGPT’s outputs within real-world team workflows. Our findings reveal that developers frequently use ChatGPT for purposes beyond code generation, such as conceptual guidance, naming, documentation refinement, and debugging support. We observe a spectrum of use, from direct integration to complete rejection, shaped by contextual judgment, trust, and project norms. This perspective contributes a more grounded understanding of AI-assisted collaboration, informing both tool design and responsible adoption.
The paper makes the following contributions:
We introduce and apply PatchTrack, a classifier for identifying how developers integrate, adapt, or reject ChatGPT-generated code in real pull requests.
We provide a mixed-methods analysis of applied patches, identifying recurring patterns of structural integration, selective extraction, and iterative refinement, showing that developers treat AI-generated code as a starting point rather than a final solution.
We show that ChatGPT influences pull request outcomes even when no code is adopted, through conceptual guidance, documentation refinement, and debugging support.
We identify cases where valid AI-generated patches are discarded due to architectural misfit, contributor role, or workflow norms, highlighting socio-technical barriers to adoption.
@@
We release a public replication package [42] to support reproducibility and future research on AI-assisted collaboration.
@@ 2.1 Terminology
PatchTrack: A classification tool that determines whether ChatGPT-generated patches were applied, not applied, or not suggested during developers’ integration with ChatGPT.
@@ 2.2 Classifying Developer Interaction with ChatGPT
A typical GitHub pull request workflow augmented with ChatGPT interactions can originate from the pull request author or a reviewer. Consider a scenario in which a developer, intending to fix a bug, add a feature, or improve an existing one, initiates a change to a file named Foo. Following a pull-based development workflow [15], the developer creates a new branch, commits changes, and eventually merges the pull request into the main repository of the project.
@@
If the code snippets are not applied, or modified beyond recognition (resulting in a semantic clone), then Foo may remain unchanged or include changes influenced by the discussed concepts. This scenario is defined as “Patch Not Applied (PN)”.
@@
This categorization underpins our analysis by structuring the diverse ways developers interact with ChatGPT in pull requests. By distinguishing between directly applied patches (PA), substantially modified or conceptually reused suggestions (PN), conceptual-only interactions without code (NE), and rejected contributions in closed pull requests (CL), we can (i) trace ChatGPT’s concrete influence on code changes, (ii) identify barriers to adoption, trust, and contextual fit, (iii) examine decision-making dynamics in AI-assisted workflows, and (iv) analyze failure cases where AI involvement did not lead to integration. The inclusion of CL cases is particularly valuable for understanding how project constraints, scope misalignment, or collaboration breakdowns affect AI-supported contributions. This classification also informs targeted recommendations: improving the usability of the patch (PA), aligning suggestions with developer expectations (PN), improving conceptual guidance (NE) and mitigating risks in unmerged PRs (CL). Together, these categories link observed outcomes to distinct usage patterns, supporting evidence-based insights into how generative AI can enhance the efficiency and reliability of collaborative code review.
@@ 3.1 Research Questions
RQ1: How can PatchTrack classify the use and integration of ChatGPT-generated patches in pull requests under self-admitted ChatGPT usage? This RQ investigates both the distribution and nature of ChatGPT-generated patch integration in merged pull requests. We use PatchTrack to classify pull requests into three categories: (i) patches that were applied (PA), (ii) patches that were suggested but not applied (PN), and (iii) cases where ChatGPT did not generate patches (NE). In addition to this quantitative view, we conduct a thematic analysis of PA cases to understand how developers adopt, reshape, or refine ChatGPT’s contributions. Together, these perspectives provide a comprehensive view of AI-generated patch adoption and its role in collaborative development.
@@
3.1.1 Scope and Unmodeled Factors.
Our study does not control for prompt formulation or developer expertise, though both can influence the quality and integration of AI-generated code. Prior work shows that incomplete or ambiguous prompts can yield suboptimal suggestions [46], and that developers often adapt or reject AI-generated code based on perceived quality [10, 11]. We instead focus on self-admitted ChatGPT usage in real-world pull requests, enabling grounded analysis of integration behavior without assumptions about prompt content. These unmodeled factors affect all classes in our study (PA, PN, NE, CL) and are unlikely to bias classification or thematic coding disproportionately. We highlight prompt quality and user expertise as valuable future directions for modeling variation and supporting cross-tool comparisons in AI-assisted development.
3.1.2 Rationale for Categorization
We classify ChatGPT-assisted pull requests into four categories: Patch Applied (PA), Patch Not Applied (PN), No Patch Generated (NE), and Closed (CL). This categorization is central to our study as it enables a structured understanding of how developers engage with AI-generated suggestions. PA cases allow us to trace direct integration of ChatGPT-generated code. PN cases highlight friction points such as lack of trust, scope misalignment, or manual refinement that prevent full adoption. NE cases reveal how developers use ChatGPT for conceptual understanding or planning, even when no code is generated. Finally, CL cases expose collaborative or technical breakdowns where AI-supported contributions did not result in a merged PR. By segmenting our dataset in this way, we surface distinct usage patterns, support comparative analysis across outcomes, and frame actionable insights for improving AI-assisted pull request workflows.
@@ 3.2.1 Step 1: Identification of merged pull requests containing ChatGPT conversation shared links:
This process identified 187 new merged PRs and 14 closed/unmerged PRs. We excluded only PRs whose ChatGPT share links were inactive (e.g., deleted or inaccessible). We did not filter repositories by popularity or overall activity, but required that each repository contain at least one PR with at least one review comment to ensure meaningful discussion context. The final dataset comprises 338 PRs from 255 unique repositories. Summary repository metrics (median stars = 58, median forks = 31, median watchers = 946) indicate that most projects are active and nontrivial; full distributions are presented in Section 4.1 (Fig. 3).
@@ 3.2.2 Step 2: Extracting pull requests and ChatGPT patches
The pull request and ChatGPT patches are central for analysis in RQ1. Our dataset consists of links to pull requests, ChatGPT links, and patches (if any) extracted from the conversation. The process of extracting patches from ChatGPT conversations was not completely automated, especially in the extended dataset. We noted that OpenAI’s updated terms of service (as of November 14, 2023) prohibit users from automatically extracting data directly from their services [48]. For this reason, we did not employ any automation or scraping tools on ChatGPT’s website or API. Instead, we manually converted each shared conversation link obtained in Step 1 into a local HTML file and performed any necessary parsing only on these static offline copies local without making HTTP requests or automated interaction with OpenAI’s systems. Pull request patches were extracted using GitHub’s /pulls/ endpoint and saved as “diff” files. At the end of this step, we had two sets of patches for each pull request: ChatGPT and pull request patches. This data served as input for the patch classification phase.
@@ 3.2.3 Step 3: Patch classification
To answer RQ1, we needed a reliable way to determine whether and how ChatGPT-suggested patches were integrated into pull requests. This required automatically identifying ChatGPT code snippets and comparing them against the patches actually merged by developers. The goal of this step is to classify each pull request into one of several categories, such as Patch Applied (PA), Patch Not Applied (PN), or No Patch Exists (NE), based on how the AI-generated patch relates to the developer’s final code.
To enable this classification at scale, we developed PatchTrack, an automated tool that analyzes the presence, absence, or partial use of ChatGPT-suggested code in merged pull requests. The input to this step consists of (i) ChatGPT conversation links extracted from pull requests and (ii) the corresponding set of code diffs (patch hunks) introduced in the pull request. The output is a patch-level label indicating whether the ChatGPT code was directly applied, discarded, or not present in the conversation. This automated labeling supports large-scale, post-hoc analysis of developer behavior under self-admitted ChatGPT usage, which is difficult to assess via metadata or manual inspection alone.
@@
Building on this foundation, we adapted PaReco [49] to create PatchTrack, tailoring it to detect and classify ChatGPT-influenced patches in pull requests. While PaReco was designed for the detection of code clones within divergent variant forks, identifying missed opportunities, effort duplication, and split cases between mainline and variant branches, it is not directly applicable for detecting ChatGPT patches. In contrast, PatchTrack specifically identifies ChatGPT-shared links, extracts patches, and analyzes prompts. Although PaReco supports exact token matching, its reliance on a Bloom filter data structure leads to higher false positive rates, which can undermine patch identification precision. PatchTrack addresses this limitation by using a temporary array of hash tables for token matching, allowing precise matches with reduced false positives. Hash tables are particularly suitable for exact data retrieval, offering an average-case time complexity of for lookups [50, 51], which contrasts favorably with the lookup time associated with Bloom filters. This design significantly improves PatchTrack’s efficiency and accuracy, making it more effective for exact token matching in diverse software development contexts.
To support robust patch comparison, PatchTrack applies normalization to both ChatGPT and PR code snippets. This includes removing whitespace and comments, converting to lowercase, filtering non-ASCII characters, and applying syntax-aware rules based on file type. The tool currently supports 38 file types, covering all major languages in our dataset, and applies tailored heuristics to account for syntax and comment style differences. Its modular design allows scalable extension to additional formats, ensuring robust and adaptable classification across diverse codebases.
@@
A patch is classified as PA, PN, NE, “cannot classify” (CC), or “error encountered” (EE). This classification occurs at the hunk level within each pull request. For each hunk, PatchTrack compares the n-gram tokens of the added lines (prefixed with ‘’) to the tokens in the ChatGPT snippet. If at least one match is found, it is labeled PA; otherwise, it is PN. If no ChatGPT code snippet exists in the conversation, all hunks are labeled NE. Unsupported file types result in CC; processing failures are labeled EE.
To determine the optimal n-gram size, we evaluated the classification accuracy of PatchTrack under . Table 1 shows the distribution of classified PRs across PA, PN, and NE categories for each configuration. Table 2 presents performance metrics, accuracy, precision, recall, and F1 score, based on manual validation of a sample of 230 cases.
| Classification Type | n = 1 | n = 2 | n = 3 | n = 4 |
| Patch Applied (PA) | 116 (40.7%) | 83 (29.1%) | 47 (16.5%) | 38 (13.3%) |
| Patch Not Applied (PN) | 63 (22.1%) | 96 (33.7%) | 132 (46.3%) | 141 (49.5%) |
| No Patch Exists (NE) | 106 (37.2%) | 106 (37.2%) | 106 (37.2%) | 106 (37.2%) |
| N-Gram | Class | Acc. | Prec. | Recall | F1 |
| n = 1 | PA | 97.2 | 96.6 | 95.6 | 96.1 |
| PN | 95.4 | 92.9 | 94.5 | 93.7 | |
| NE | 100 | 100 | 100 | 100 | |
| Overall | 97.5 | 96.5 | 96.7 | 96.6 | |
| n = 2 | PA | 95.6 | 95.3 | 88.4 | 91.7 |
| PN | 95.3 | 90.3 | 96.2 | 93.2 | |
| NE | 100 | 100 | 100 | 100 | |
| Overall | 96.9 | 95.2 | 94.7 | 94.9 | |
| n = 3 | PA | 95.9 | 88.1 | 88.1 | 88.1 |
| PN | 95.9 | 94.9 | 94.9 | 94.9 | |
| NE | 100 | 100 | 100 | 100 | |
| Overall | 97.2 | 94.3 | 94.3 | 94.3 | |
| n = 4 | PA | 96.7 | 90.9 | 85.7 | 88.2 |
| PN | 94.5 | 95.2 | 97.5 | 96.3 | |
| NE | 100 | 100 | 100 | 100 | |
| Overall | 97.1 | 95.3 | 94.4 | 94.8 |
We selected as the final configuration, since it yielded the highest overall accuracy and better captured short or partial code fragments common in ChatGPT responses. Larger values led to reduced recall and more PN classifications due to stricter token match requirements.
To validate this choice, we randomly sampled 230 pull requests (90 PA, 55 PN, 84 NE) from the full set of 285 PRs classified by PatchTrack (116 PA, 63 PN, 106 NE), using a 95% confidence level and 5% margin of error, and compared PatchTrack’s output to human labels. Each ChatGPT-generated patch was matched against all patches in the pull request using consistent labeling criteria. Inter-rater agreement between the two authors yielded a Cohen’s Kappa of 0.85, indicating strong reliability. Performance metrics across different n-gram sizes are shown in Table 2, and class distributions are shown in Table 1. Based on this comparative analysis, we selected as the optimal configuration, as it achieved the best balance of accuracy, precision, and recall while capturing short, syntactically varied code fragments typical of ChatGPT output. To better understand PatchTrack’s limitations, we manually inspected 7 misclassified cases under the setting: 4 cases labeled PA that should be PN, and 3 labeled PN that should be PA. Two main error sources emerged:
GitHub API Limitations: GitHub omits diffs larger than 400 lines in its API responses. Three misclassifications stemmed from ChatGPT code being present in these omitted sections, leading to false PN results.
Single-Line Matches: In some cases, short one-line overlaps triggered PA labels, even when they lacked semantic relevance. These borderline matches reveal a trade-off between sensitivity and precision.
While PatchTrack performs well for direct matches, it does not capture semantic or structural adaptations of AI-generated code. Future extensions may incorporate AST or embedding-based approaches to address this limitation.